# Imitation Game: Reproducing Deep Learning Bugs Leveraging an Intelligent Agent

### v1

@@ 4.6.4. Answering RQ4 – Usefulness
To evaluate the usefulness of our generated code in a practical setting, we conducted a controlled study involving 27 professional developers and AI engineers (n=14 experimental, n=13 control). To ensure that our developer study had sufficient statistical sensitivity capable of detecting meaningful differences between two experimental settings, we conducted an a priori power analysis. This analysis estimates whether the planned sample size can reliably detect effects of practical significance. The results indicate that our study is adequately powered (80% power, $\alpha$=0.05) to detect large effects (Cliff’s $\delta\geq$ 0.47) in continuous measures (e.g., time) and substantial differences in success rates. This aligns with typical effect sizes observed in the software engineering studies for tool evaluation and is appropriate for detecting meaningful practical improvements in bug reproduction (Kampenes et al., 2007; de Oliveira Neto et al., 2019). To analyze the results, we employed the same strategy as in prior work (Shah et al., 2025).
@@
We also assessed the cognitive load of the participants using the NASA Task Load Index (NASA-TLX). As shown in Fig. 6, the experimental group reported significantly lower cognitive burden across all dimensions, including a 43.2% $\pm$ 28.2% reduction in frustration and a 25.6% $\pm$ 23.1% reduction in mental demand. Furthermore, participants in the experimental group also reported higher performance (4.21 vs. 3.23), corresponding to a 30.34% $\pm$ 24.46% improvement relative to the control group. Based on our collected responses, the experimental group also rated the overall usefulness of our generated code at 4.57 out of 5, with 92.30% rating it as ”Useful” or ”Highly Useful” (¿4). All participants in the experimental group would also prefer the inclusion of the generated code in bug reports. Thus, all the findings above suggest that RepGen’s generated code offers significant benefits by improving bug-reproduction success rates, reducing time spent, and lowering developers’ cognitive load.

### v2

@@ 4.6.4. Answering RQ4 – Usefulness
To evaluate the usefulness of our generated code in a practical setting, we conducted a controlled study involving 27 professional developers and AI engineers (n=14 experimental, n=13 control). To ensure that our developer study had sufficient statistical power to detect meaningful differences between two experimental settings, we conducted an a priori power analysis. This analysis estimates whether the planned sample size can reliably detect effects of practical significance. The results indicate that our study is adequately powered (80% power, $\alpha$=0.05) to detect large effects (Cliff’s $\delta\geq$ 0.47) in continuous measures (e.g., time) and substantial differences in success rates. This aligns with typical effect sizes observed in software engineering studies on tool evaluation and is appropriate for detecting meaningful, practical improvements in bug reproduction (Kampenes et al., 2007; de Oliveira Neto et al., 2019). To analyze the results, we employed the same strategy as in prior work (Shah et al., 2025).
@@
We also assessed participants’ cognitive load using the NASA Task Load Index (NASA-TLX). As shown in Fig. 6, the experimental group reported significantly lower cognitive burden across all dimensions, including a 43.2% $\pm$ 28.2% reduction in frustration and a 25.6% $\pm$ 23.1% reduction in mental demand. Furthermore, participants in the experimental group also reported higher performance (4.21 vs. 3.23), corresponding to a 30.34% $\pm$ 24.46% improvement relative to the control group. Based on our collected responses, the experimental group also rated the overall usefulness of our generated code at 4.57 out of 5, with 92.30% rating it as ”Useful” or ”Highly Useful” (¿4). All participants in the experimental group would also prefer the inclusion of the generated code in bug reports. Thus, all the findings above suggest that RepGen’s generated code offers significant benefits by improving bug-reproduction success rates, reducing time spent, and lowering developers’ cognitive load.
@@ 7. Conclusion & Future Work
Acknowledgments:
This work was supported by the Natural Sciences and Engineering Research Council of Canada (Discovery Grant RGPIN-03236), the Fonds de recherche du Québec (FRQ), and the Canadian Institute for Advanced Research (CIFAR).